The Rachel Company has sales of $760,000, and the break-even point in sales dollars is $638,400. Determine the company's margin of safety as a percent of current sales.
Answer #1

margin of safety=Total sales-Breakeven sales

=(760,000-638,400)=121600

margin of safety%=margin of safety/Total Sales

=121600/760,000

=16%